The mean height of random sample of 100 individuals from a population is 160 cms. The


S.D. of the sample is 10cms. Would it be reasonable to suppose that the mean height of the


population is 165cms?

1
Expert's answer
2022-05-26T06:26:09-0400

Given that "n=100, \=x=160, \sigma=10, \mu=165"

"H_0:\mu=165" (there is no difference between sample mean and opulation mean)

"H_a: \mu\not=165" (two-tailed alternative)

The test statistic is given by

"z=\frac{\=x-\mu}{\sigma/\sqrt{n}}\\=\frac{160-165}{10/\sqrt{100}}=-5"

At 5% significance level the tabulated value for "z_{\alpha}=1.96"

|Calculated value|"\le" Tabulated value then Accept "H_0"

"|5|>1.96" So we reject "H_0"

That is there is a significant difference between the sample mean and population means.